OpenStudy (anonymous):

Determine if the function is either even, odd or neither: x^2-x^8+1

OpenStudy (anonymous):

if f(x)=f(-x) then f is even if f(x)=-f(-x) then f is odd now u have f(x)=x^2-x^8+1 and f(-x)=x^2-x^8+1 then what is result?

OpenStudy (anonymous):

even?

OpenStudy (anonymous):

or does the plus one make it odd?

OpenStudy (anonymous):

u are right 'even'

OpenStudy (anonymous):

okay so the plus one has no effect

OpenStudy (anonymous):

no effect

OpenStudy (anonymous):

okay i have one more question

OpenStudy (anonymous):

 this represents an odd function correct?